# Shortcut to Chemically Accurate Quantum Computing via Density-based Basis-set Correction

### v1

@@ II.1 Ground-state Energies
Ground-state energies are presented in Figure 2 and Table 2. Details regarding the ADAPT-VQE iterations and the associated “ADAPT” values are provided in Fig. 1 to 14 in the SI, as long as additional tests on hydrogen chains and atoms. In Figure 2, a general trend is observed where the basis-set corrected energies align with value between the cc-pVDZ and cc-pVTZ levels, involving fewer than twenty qubits. As we can see from Table 2, both strategies provide the same quantitative improvements. This shows that the results are consistent with the conclusion from Ref.[12] where the authors justified Strategy 1. We remind here that the basis-set correction does not fix the intrinsic error of the quantum algorithm.
We applied the basis-set correction strategies to systems with different correlation regimes, and for the best cases, we reach errors in the order of tens of mHa: 40 mHa for H2O with 24 qubits, 60 mHa for N2 with 16 qubits, and less than ten mHa errors for LiH and H2 with less than 10 qubits. Without the basis-set correction, these results would have required more than a hundred of logical qubits. The results stay consistent for the other systems available in the SI (H4, H6, H8). For H2 using the cc-pVDZ basis set, the optimization reached chemical accuracy relative to the CBS limit, a process that used 20 logical qubits. Similar convergence to the CBS limit is observed for He and Be (see SI). Notably, this outcome contrasts with scenarios that typically require more than two hundred logical qubits. These results reflect findings from previous publications on the DBBSC method[8], where significant corrections typically begin with at least the cc-pVTZ basis set. Consequently, if one had access to a few hundred logical qubits, achieving chemical accuracy would be theoretically feasible for all presented cases. To support this claim, we have reported classically computed values for CIPSI energy corrections using Dunning basis sets (see Table 1 of the SI). According to this Table, the error relative to the CBS energy for the LiH molecule is 0.2 mHa with the cc-pVTZ basis set and both the correlation and HF basis-set corrections. For H2O, we achieve an error of 3 mHa from the cc-pVTZ basis set, and for N2, the error is 1 mHa from the cc-pVTZ basis set. These results suggest that chemically relevant accuracies are achievable with around a hundred of logical qubits. However, the focus of our applications extends beyond simple energy calculations; we also examine energy derivatives through the dissociation curve, which will be discussed in the following part.
Since the discussed qubit counts were obtained using standard basis sets, we report in Table 2 computations using selected SABS basis sets for LiH, N2, and H2O. They were chosen to match the sizes of basis sets used in previous computations. Specifically for H2O, employing the cc-pV5Z-10 basis (24 qubits) achieves results slightly superior to those obtained with the 6-31G basis set, using the same number of iterations. For H2O, we observed a reduction in the HF correction by 40 mHa when moving from the 6-31G to the cc-pV5Z-10 basis, by over 100 mHa when comparing the STO-3G to the cc-pV5Z-4 basis for the LiH molecule, and more than an Hartree for the nitrogen molecule moving from the STO-3G to the cc-pV5Z-6 basis. Based on these findings, we anticipate that further exploration of this strategy could lead to a correction scheme requiring a single correction term. Of course, the interest of SABS is to be systematically improvable in the direction to the full accuracy of the corresponding Dunning basis set. It is important to note that this SABS systematic improvement is not limited to quantum algorithms such as ADAPT-VQE and is also present for the HF and FCI energies. Since the SABS approach strongly reduces the computational cost while maintaining accuracy, it should offer further applications of the CIPSI approach in classical quantum chemistry, especially when using the basis-set HF correction using large basis-set HF computations. In practice, a cc-pVTZ-like quality is reached (see Table 2) for N2 with a cc-pV5Z-11 basis set (32 qubits) and for H2O with a cc-pV5Z-11 (30 qubits) basis set. A cc-pV5Z-10 basis set (28 qubits) achieves a cc-PV5Z-like ground-state energy accuracy for LiH. Overall, the SABS always provide the best SC-ADAPT solutions (i.e. before the HF correction). Finally, one last aspect of the analysis of such large basis-set simulations is related to the convergence of ADAPT-VQE. Indeed, in VQE-type computations [5], there is no formal guarantee of convergence and ADAPT-VQE belong to such heuristic family of methods. If, for systems like H2, LiH or N2 (see SI, Figures 3-14), convergence is obtained in a few dozen of iterations, for more complex systems such as H2O, the number of required iterations strongly increases when a double-zeta basis set is used. Clearly, Figure 2 in the SI shows that more than a thousand iterations would be required to achieve full convergence starting from the HF initial state. This heuristic aspect makes that no anticipation of the exact required number of iterations can be made. Moreover, if this could be fine on a real quantum processor, the use of classical emulation makes each iteration more costly in term of time-to-solution than the previous, limiting the overall convergence capabilities. It is possible to force convergence by replacing the HF starting point by a CIPSI one. Figure 2 in the SI shows that chemical accuracy can be easily reached with a well-converged CIPSI solution. However, ADAPT-VQE struggles to improve the solution which is expected due to the enormous size of the parameter space. Furthermore, the CIPSI-based initial state already contains significant amount of correlation. This represents a hard challenge for the ADAPT-VQE procedure, which now needs to more carefully pick the next ansatz operator to improve the existing quantum state. In any case, starting from an unconverged CIPSI wave function is a robust solution to strongly reduce the overall computational time. One example is given for H2O and the V5Z-10 basis set which initially led to a not fully converged result. With a better CIPSI starting point, it is possible to recover a few milli-Hartrees (see numbers in parenthesis in Tab. 2). It is also possible to change the operator pool but, the point here is that convergence becomes challenging when tackling complex electronic structures and such computations would not be possible without GPU-accelerated emulation. Overall, concerning the basis sets, it is important to point out a key anomaly, i.e. the remarkable performance of the minimal STO-3G basis set. This was expected as Pople already discussed the outperformance of this basis set in the seventies [40]. Indeed, Davidson and Feller detailed in their 1986 review [41] the existence of error compensations and stated that “the smaller the basis the more ab initio calculations assume an empirical flavor”. In practice, among the available minimal basis possibilities, STO-3G proved to be the most robust and cost-efficient choice for the DBBSC basis-set corrections. Also, in the DBBSC schemes, the STO-3G basis set highly benefits from the HF basis-set correction which reduces as the basis set and number of qubits increases. When compared to SABS, for example, STO-3G always display smaller SC-ADAPT values and larger HF basis-set corrections. In practice, Pople also noted that STO-3G is especially good for energies around the minimum, and larger basis sets are clearly required to describe accurately the notoriously more difficult dissociation curves and dipole moments.
@@
Table 2: Ground-state energies (in Hartree) for H2O, N2, LiH, H2, and H8 calculated by FCI, ADAPT-VQE (denoted ADAPT), and basis-set corrected ADAPT-VQE according to Strategy 1 (denoted as A+$\Delta$c+$\Delta$HF) and to Strategy 2 (denoted as SC-A and SC-A+$\Delta$HF). $\Delta$c corresponds to the non-self-consistent correlation basis-set correction and $\Delta$HF corresponds to the HF basis-set correction. The frozen-core approximation has been used for H2O, N2, and LiH. The CBS limits are estimated by two-point extrapolations from cc-pVQZ and cc-pV5Z calculations. (1) Value between parentheses corresponds to a CI starting points with 11016 determinants selected with the CIPSI procedure.
[…]
@@
pcseg-0 24 -75.90855 -75.90843111500 iter. with 14668 CI selected with the CIPSI procedure for the initial state. -76.33681 -76.03999 -76.33279
6-31G 24 -76.11995 -76.119892221000 iter. with 10879 CI for the initial state. -76.28035 -76.23717 -76.32025
V5Z-10 24 -76.12626 -76.12409 -76.32705 -76.27418 (.27636 )(1) -76.32505
@@
Table 3: Dipole moments of H2O and LiH, frozen core calculations, FCI, and self-consistently basis-set corrected ADAPT-VQE (noted SC-A) in atomic units. $\Delta$HF corresponds to the HF basis-set correction to the dipole moment. We use the difference between the aug-cc-pV5Z Hartree-Fock dipole moment with the Hartree-Fock dipole moment in the basis of interest to compute $\Delta$HF. Details are available in Table 5 of the SI.
[…]
@@
Figure 3: Dissociation curves of H2, LiH and N2 molecules. Blue and purple lines correspond to Dunning basis sets. Red and green plain lines correspond to ADAPT-VQE calculations without basis-set corrections. Red and green dash-dotted lines correspond to basis-set corrected dissociation curves.
II.2 Dissociation Curves
We pursue with the dissociation energies reported in Figure 3. Clearly, the simple ADAPT-VQE/STO-3G level of theory (see Figure 3 and Table II) appears quite far from an accurate description of the systems dissociation (i.e. compared to the largest basis sets), which starts with FCI/triple-zeta (and beyond) levels. The non-corrected ADAPT-VQE values are represented with bold lines whereas the corrected values are reported with dashed curves. These dissociation curves are extremely challenging as they require to capture both electronic correlation and HF basis-set corrections. First, we notice that for all the cases, the corrected values around the equilibrium are always substantially closer to the large cc-pV5Z reference basis-set values than the uncorrected ones and improve with the basis-set size and the use of SABS. For longer-range distances, the basis-set requirements appear even more stringent. In practice, the 3 dissociation curves manage to converge to high accuracy thanks to the use of SABS: nearly a triple-zeta quality for N2 (using a corrected cc-pV5Z-10 basis), nearly to a cc-pV5Z quality for H2 (using a corrected cc-pV5Z-8 basis) and nearly a cc-pV5Z for LiH (using a corrected cc-pV5Z-8 basis). For H2, the corrected cc-pVDZ curve matches the cc-pV5Z reference until 2.5 Angstrom and a slight discrepancy appears at long range. This is consistent with classical computations leading to a CBS convergence with a corrected cc-pVTZ basis (114 qubits). It is possible to fix this issue more affordably by enriching ADAPT-VQE using a cc-pV5Z-8 basis requiring only 24 qubits. For LiH, it is also possible to fix the small discrepancies observed on the cc-V5Z-7 basis-set dissociation curve using a larger cc-V5Z-10 (28 qubits) basis which converges to the CBS limit. Finally, it is important to highlight the good DBBSC performance for the prediction of the triple-bonded N2 dissociation curve. Indeed, such computation is well-documented in the literature and known as extremely difficult as it requires both a multi-reference treatment and core-valence basis sets [42, 43, 44]. At the cost of a minimal basis set (i.e. 16 qubits), our N2 DBBSC computations achieved nearly a triple-zeta (cc-pVTZ) quality providing an accuracy which would have required around 100 logical qubits in the context of a brute-force simulation. To the best of our knowledge, these results are the most accurate using a quantum algorithm when compared to the most recent results from the literature. Indeed, a group of researchers managed to predict such dissociation using a Local Unitary Cluster Jastrow (LUCJ) ansatz coupled to double-zeta basis sets (6-31G and cc-pVDZ) [45]. The computation required the use of massive computational resources, namely hundreds of compute nodes of the Fugaku classical supercomputer coupled to a QPU. Alternatively, in the context of the fermionic quantum emulator, N2 simulations used ADAPT-VQE coupled to the 6-31G and def2-SVP (56 qubits) basis sets to perform computation via approximate Matrix Product States (MPS) [46]. We achieved a better accuracy with far less qubits and, again, these results can be systematically improved using larger SABS. Indeed, going beyond a cc-pVTZ basis set is possible using the next (larger) SABS in term of size: cc-pV5Z-11 (32 qubits).
@@ II.3 Dipole Moments
We also explored the idea to extend the basis-set correction schemes to molecular properties such as the dipole moment. We report in Table 3 the dipole moments of H2O and LiH. In practice, it is known that dipole moments require the addition of diffuse functions for which the DBBSC approximation does not correct. However, according to Ref. [47], the rate of convergence of the dipole moment with the basis size should be the same as for the energy. Therefore, it is relevant to apply the basis-set correction to the dipole moment. Here, the chemical accuracy refers to the best estimation provided using the cc-pV5Z basis set. They are computed in the classical software from the density out of the last ADAPT-VQE run. From Table 3, the second strategy leads to dipole moments with errors slightly superior to FCI dipole moments. However, the addition of the HF basis-set correction permits to reach dipole moments comparable to the double-zeta calculations for both systems. It is important to note that the observed ADAPT-VQE slow convergence in term of iterations is the source of the small errors. For LiH, the STO-3G dipole moment reaches an error of 0.01 atomic units. For H2O, the error is reduced to 0.01 atomic units with respect to the cc-pVDZ value which has an error of 0.2 atomic units. In all cases, strong improvements are observed with respect to the STO-3G/ADAPT-VQE level. Further improvements are observed as the size of the basis sets increases.

### v2

@@ II.1 Ground-state Energies
Ground-state energies of the H2, LiH, H2O, and N2 molecules close to their respective equilibrium geometries are presented in Figure 2 and Table 2. Details regarding the ADAPT-VQE iterations and the associated “ADAPT” values are provided in the SI, as well as additional tests on atoms and hydrogen chains.
@@
In Figure 2, a general trend is observed: the basis-set corrected ground-state energies with the small basis sets, i.e. STO-3G, 6-31G, or pcseg-0 [43], align with values between the cc-pVDZ and cc-pVTZ basis-set levels, while requiring much less qubits than these latter basis sets. As we can see from Table 2, both strategies provide the same quantitative improvements. These results are consistent with the conclusion from Ref. [13] where the non-self-consistent approximation was found sufficient for calculating energies.
For the best cases, the basis-set corrected ground-state energies have errors in the order of tens of mHa: 40 mHa for H2O with 24 qubits, 60 mHa for N2 with 16 qubits, and less than 10 mHa for LiH and H2 with less than 10 qubits. The results stay consistent for the other systems available in the SI (H4, H6, H8). For H2 using the cc-pVDZ basis set, the basis-set corrected ground-state energy reaches chemical accuracy relative to the CBS limit, while requiring 20 logical qubits. Similar convergence to the CBS limit is observed in the SI for He and Be. Without the basis-set correction, reaching a similar accuracy would have required typically more than a hundred of qubits.
In the initial work on the DBBSC method [7], it was found that basis-set corrected ground-state energies of atoms reach chemical accuracy with the cc-pVTZ basis set. Similarly, for the molecules studied here, if one had access to a few hundred logical qubits required for the cc-pVTZ basis set, achieving chemical accuracy would be theoretically feasible for all cases. To support this claim, we report in the SI classically computed basis-set corrected FCI ground-state energies using Dunning basis sets. The error relative to the CBS limit for the LiH molecule is 0.2 mHa with the cc-pVTZ basis set. For H2O and N2, we achieve errors of 3 mHa and 1 mHa, respectively, with the cc-pVTZ basis set.
@@
Table 2: Ground-state energies (in Ha) for H2O, N2, LiH, H2, and H8 calculated by FCI, ADAPT-VQE (denoted as ADAPT), and basis-set corrected ADAPT-VQE according to Strategy 1, denoted as A+PBE+$\Delta$HF, and to Strategy 2, denoted as SC(A+PBE) and SC(A+PBE)+$\Delta$HF (i.e., without and with the HF basis-set correction, respectively). Here, PBE refers to the PBE-based correlation basis-set correction and $\Delta$HF refers to the HF basis-set correction, and SC stands for “self-consistent”. The frozen-core approximation has been used for H2O, N2, and LiH. The CBS limits are estimated by two-point extrapolations from cc-pVQZ and cc-pV5Z calculations.
H2O $N_{\text{qubits}}$ FCI ADAPT A+PBE+$\Delta$HF SC(A+PBE) SC(A+PBE)+$\Delta$HF
@@
pcseg-0 24 -75.90855 -75.90843111500 iterations with a 14668-determinant CIPSI initial state. -76.33681 -76.03999 -76.33279
6-31G 24 -76.11995 -76.119892221000 iterations with a 10879-determinant CIPSI initial state. -76.28035 -76.23717 -76.32025
V5Z-10 24 -76.12626 -76.12409 -76.32705 -76.27418333The value is -76.27636 when using a 11016-determinant CIPSI initial state. -76.32505
@@
Let us now discuss the results obtained with our SABS for LiH, N2, and H2O. They were chosen to match the sizes of the small basis sets previously discussed. Specifically for H2O, employing the V5Z-10 basis set (24 qubits) achieves results slightly superior to those obtained with the 6-31G basis set. For H2O we observed a reduction in the HF basis-set correction by 40 mHa when moving from the 6-31G to the V5Z-10 basis set, by over 100 mHa when comparing the STO-3G to the V5Z-4 basis set for the LiH molecule, and more than an Hartree for the N2 molecule moving from the STO-3G to the V5Z-6 basis set. Based on these findings, we anticipate that further exploration of this strategy could lead to a basis-set correction scheme requiring only the correlation basis-set correction term. Of course, the interest of SABS is to be systematically improvable toward the parent Dunning basis set. It is important to note that this SABS systematic improvement is not limited to quantum algorithms such as ADAPT-VQE and is also present for the HF and FCI calculations. Since the SABS approach strongly reduces the computational cost while maintaining accuracy, it should offer further applications of the CIPSI approach in classical quantum chemistry. In practice, a cc-pVTZ-like quality is reached for N2 with a V5Z-11 basis set (32 qubits) and for H2O with a V5Z-11 basis set (30 qubits). A V5Z-10 basis set (28 qubits) achieves a cc-pV5Z-like ground-state energy accuracy for LiH. Overall, the SABS always provide the best ADAPT energies with the self-consistent correlation basis-set correction (before addition of the HF basis-set correction).
Overall, concerning the basis sets, it is important to point out a key anomaly, i.e. the remarkable performance of the minimal STO-3G basis set. This was expected as Pople already discussed the outperformance of this basis set in the seventies [44]. Indeed, Davidson and Feller detailed in their 1986 review [45] the existence of error compensations and stated that “the smaller the basis the more ab initio calculations assume an empirical flavor”. In practice, among the available minimal basis-set possibilities, STO-3G proved to be the most robust and cost-efficient choice for the basis-set corrections. Also, in the DBBSC schemes, the STO-3G basis set highly benefits from the HF basis-set correction which reduces as the basis set and number of qubits increases. When compared to SABS, for example, STO-3G always display smaller self-consistent basis-set corrected ADAPT values and larger HF basis-set corrections. In practice, Pople also noted that STO-3G is especially good for energies around the equilibrium geometry, and larger basis sets are clearly required to describe accurately the notoriously more difficult dissociation curves and dipole moments.
Finally, one last aspect of the analysis of such large basis-set simulations is related to the convergence of ADAPT-VQE. Indeed, in VQE-type computations [5], there is no formal guarantee of convergence and ADAPT-VQE belongs to such heuristic family of methods. If, for systems like H2, LiH or N2 (see SI, Figures 3-14), convergence is obtained in a few dozen of iterations, for more complex systems such as H2O, the number of required iterations strongly increases when a double-zeta basis set is used. Clearly, Figure 2 in the SI shows that more than a thousand iterations is required to achieve full convergence starting from the HF initial state. This heuristic aspect makes that no anticipation of the exact required number of iterations can be made. This may be manageable on a real quantum processor, but the use of classical emulation makes each iteration more costly in term of time-to-solution than the previous one, limiting the overall convergence capabilities. It is possible to force convergence by replacing the HF starting point by a CIPSI one. Figure 2 in the SI shows that chemical accuracy can be easily reached with a well-converged CIPSI solution. However, ADAPT-VQE struggles to improve the solution which is expected due to the enormous size of the parameter space. Furthermore, the CIPSI-based initial state already contains significant amount of correlation. This represents a hard challenge for the ADAPT-VQE procedure, which now needs to more carefully pick the next ansatz operator to improve the existing quantum state. In any case, starting from an unconverged CIPSI wave function is a robust solution to strongly reduce the overall computational time. One example is given for H2O and the V5Z-10 basis set which initially led to a not fully converged result. With a better CIPSI starting point, it is possible to recover a few milli-Hartrees (see footnotes in Table 2). It is also possible to change the operator pool but the point here is that convergence becomes challenging when tackling complex electronic structures and such computations would not be possible without GPU-accelerated emulation.
Figure 3: Dissociation curves of H2, LiH and N2 molecules. Blue and orange lines correspond to Dunning basis sets using the FCI method. The notation VXZ stands for the basis set cc-pVXZ. Green and purple plain lines correspond to ADAPT-VQE calculations without basis-set corrections. Green and purple dash-dotted lines correspond to basis-set corrected dissociation curves. Raw data are available in the SI.
II.2 Dissociation Curves
We pursue with the dissociation energies reported in Figure 3. Clearly, the simple ADAPT-VQE/STO-3G level of theory appears quite far from an accurate description of the dissociation (compared to the calculations in the largest basis sets), which is achieved with FCI/triple-zeta (and beyond) levels. The non-corrected ADAPT-VQE values are represented with bold lines whereas the corrected values are reported with dashed curves. These dissociation curves are extremely challenging as they involve several regimes of correlation going from weak correlation happening typically at short distances to strong correlation effects at large distances. First, we notice that for all the cases, the basis-set corrected values around the equilibrium are always substantially closer to the large cc-pV5Z reference values than the uncorrected ones and improve with the basis-set size and the use of SABS. For larger distances, the basis-set requirements appear even more stringent. In practice, the three dissociation curves manage to be converged to high accuracy thanks to the use of SABS: we obtain nearly a triple-zeta quality for N2 using the basis-set correction with the V5Z-6 basis set, nearly a cc-pV5Z quality for H2 using the basis-set correction with the V5Z-8 basis set, and nearly a cc-pV5Z quality for LiH using the basis-set correction with a V5Z-7 basis set. For H2, the basis-set corrected cc-pVDZ curve matches the cc-pV5Z reference up to a distance of 2.5 Å and a slight discrepancy appears at long distances. This is consistent with classical computations leading to a convergence to the CBS limit only when the basis-set correction is applied to the cc-pVTZ basis set (114 qubits). It is possible to fix this issue more affordably by using a V5Z-8 basis set requiring only 24 qubits. For LiH, it is also possible to fix the small discrepancies observed on the V5Z-7 basis-set dissociation curve using a larger V5Z-10 basis set (28 qubits) which reaches the CBS limit.
@@ II.2 Dissociation Curves
Table 3: Dipole moments (in atomic units) of LiH and H2O calculated by FCI and self-consistently basis-set corrected ADAPT-VQE, denoted as SC(A+PBE) and SC(A+PBE)+$\Delta$HF (without and with and the HF basis-set correction, respectively). Here, $\Delta$HF corresponds to the HF basis-set correction to the dipole moment, calculated as the difference between the HF dipole moment in the aug-cc-pV5Z basis set and the HF dipole moment in the considered basis set. More data are available in the SI.
LiH $N_{\text{qubits}}$ FCI SC(A+PBE) SC(A+PBE)+$\Delta$HF
@@
Finally, it is important to highlight the good performance of the DBBSC method for the prediction of the dissociation curve of the triple-bonded N2 molecule. Indeed, such computation is well-documented in the literature and known as extremely difficult as it requires both a multi-reference treatment and various weak correlation effects going from short-range to charge polarization effects [46, 47, 48, 49]. At the cost of a minimal basis set (i.e. 16 qubits), our N2 DBBSC computations achieved nearly a cc-pVTZ quality, providing an accuracy which would have required around 100 logical qubits in the context of a brute-force simulation. To the best of our knowledge, these results are the most accurate using a quantum algorithm when compared to the recent results from the literature. Indeed, a group of researchers managed to predict this dissociation curve using a Local Unitary Cluster Jastrow (LUCJ) ansatz coupled to double-zeta basis sets (6-31G and cc-pVDZ) [50]. The computation required the use of massive computational resources, namely hundreds of compute nodes of the Fugaku classical supercomputer coupled to a QPU. Alternatively, in the context of the fermionic quantum emulator, N2 simulations used ADAPT-VQE coupled to the 6-31G and def2-SVP (56 qubits) basis sets to perform computation via approximate Matrix Product States (MPS) [51]. In the present work, we achieve a better accuracy with far less qubits and, again, these results can be systematically improved using larger SABS. Indeed, going beyond a cc-pVTZ basis set is possible using the next (larger) SABS in term of size, i.e. V5Z-11 (32 qubits).
@@ II.3 Dipole Moments
We also explore the idea to extend the basis-set correction scheme in QC calculations of molecular properties such as the dipole moment. As pointed out by Halkier et al. [52], the dipole moment also suffers from a slow basis-set convergence, which can be thought of as an indirect impact of the missing short-range correlation effects in a finite basis set. Therefore, it is relevant to apply the basis-set correction to the dipole moment, as already shown in Ref. [13]. We report in Table 3 the dipole moments of H2O and LiH. We calculate the basis-set corrected dipole moment as expectation value over the dipole-moment operator over the last ADAPT-VQE wave function coming out of the self-consistent basis-set correction method. We also add a posteriori the HF basis-set correction to the dipole moment, calculated as the difference between the HF dipole moment in the aug-cc-pV5Z basis set and the HF dipole moment in the considered basis set. From Table 3, we see that the correlation basis-set correction is not sufficient to converge the dipole moments. However, the addition of the HF basis-set correction significantly improves the basis-set convergence. For both molecules, strong improvements are observed with respect to the ADAPT-VQE/STO-3G level. Further improvements are observed as the size of the basis set increases. Finally, we remark the ADAPT-VQE slow convergence in term of iterations is also the source of small errors in the dipole moments.